OpenStudy (anonymous):

if 36tt cm is the answer than the opposite is....

OpenStudy (anonymous):

I was thinking 36 cm cause that is the opposite am I right

OpenStudy (anonymous):

there is no such thing as an opposite of \(\sf 36 \pi ~cm\)

OpenStudy (anonymous):

sorry wrong question...

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

What is the radius of a circle with a circumference of 36 tt cm A.18 cm B.36 cm C.18tt cm D.72 cm -Sorry I got it mixed up "thesmartone" if u gave me a chance to write it would have been fine

OpenStudy (anonymous):

are u able to help @KAKES1967 ?

OpenStudy (anonymous):

I honestly don't know but @pooja195

OpenStudy (anonymous):

ok ..then @pooja195 - can u help

pooja195 (pooja195):

\[\huge~\rm~r=\frac{ c }{ 2\pi }\]

pooja195 (pooja195):

\[\huge~\rm~r=\frac{ 36}{ 2\pi }\]

pooja195 (pooja195):

Can you calculate the rest? Please?

OpenStudy (anonymous):

yes I was thinking 36 cm cause beore I had a question like this for 25tt cm and I got 25 cm so I say 36 cm?

pooja195 (pooja195):

I do not know where you are getting the from i want you to do the equation i gave you .

sammixboo (sammixboo):

\(\tt~r=\dfrac{ 36\pi}{ 2\pi }\)

pooja195 (pooja195):

^do that

OpenStudy (anonymous):

oh ok someonetold me that's how u do it ..I kind of figured they were wrong

pooja195 (pooja195):

@zebra8806 please dont give direct answers!

sammixboo (sammixboo):

The two \(\pi\)'s cancel each other out, so you are left with \(\tt~r=\dfrac{ 36}{ 2}\)

pooja195 (pooja195):

\[\huge~\rm~36\div2=?\]

OpenStudy (anonymous):

*facepalm* @Daddysgirl14 Instead of writing tt, please write `\(\pi\)`

pooja195 (pooja195):

^that would have helped

OpenStudy (anonymous):

18 sorry I was seeing what my mom wanted thx so its 18

pooja195 (pooja195):

Good :)

pooja195 (pooja195):

but 18 what? Pigs? Cats? Tso's?

OpenStudy (anonymous):

whoa whoa \(\huge~\rm~36\div2\neq \normalsize\text{18 sorry I was seeing what my mom wanted thx so its 18}\)

OpenStudy (anonymous):

SORRY I AM YOUNG GIRL WHATS THE PROBLEM HELPING MY PREGNANT MOM?!!

pooja195 (pooja195):

.-.

pooja195 (pooja195):

tmi

pooja195 (pooja195):

18 is right we are missing something

pooja195 (pooja195):

*but we are..

pooja195 (pooja195):

How bout a label? :P

OpenStudy (anonymous):

im not talking bout you @pooja195 im mean bout @TheЅmartOne

OpenStudy (anonymous):

lel

OpenStudy (anonymous):

what? im sorry im upset..i don't know why you are being like that thesmartone...I have pregnant mom she needs more help than I do

pooja195 (pooja195):

Are we done?

OpenStudy (anonymous):

im sorry @pooja195 im crying cause thesmartone what she said up top ^ getting mad cause I was HELPING my mom

pooja195 (pooja195):

ignore him :)

OpenStudy (anonymous):

ok ...I will also block

OpenStudy (anonymous):

:)

pooja195 (pooja195):

We can all be friends :) no need to take such actions